Find the next term in the pattern -4, 14, 10, 24,...
Found 2 solutions by mangopeeler07, stanbon:
Answer by mangopeeler07(462)

-4, 14, 10, 24,....
Find the pattern.
-4+14=10
14+10=24
10+24=34
So the next number in the pattern is 34.

Answer by stanbon(75887)

Find the next term in the pattern -4, 14, 10, 24,...
---
I ran a cubic regression program against
the data and got the following:
--------------
y = (6 2/3)x^3 -51x^2 + (124 1/3)x - 84
---
Next term: (5,6.98)
---
I doubt this is what you want but
maybe you have a better idea.
------------
Cheers,
Stan H.